# The inherent goodness of well educated intelligence

### v4

@@ I Introduction
This method of collective control takes two forms. The first, and most common, is via embedding a viscous local force on the collective. This takes the form of rules and regulations, where penalties, both financial (fines) and physical (“lock them up”), are assessed for undesirable actions. A more subtle form of viscosity was developed during the 14th and 15th century renaissance in Northern Italy lead by the Strozzi’s and Medici’s and sanctioned by the Roman Catholic Church. This financial servitude is usury or borrow/loan debt based financing (Parks, 2005). In general, economic viscosity is debt, slavery, servitude, intolerance, and lack of freedom. This is unintelligent or stupid control, where “the global consequences of the local actions” are not considered. The problem with this method of control has been that it has led to a degradation of the collective’s potential activity and an exploitative minimization of activity where fascism and capitalistic economic monopolies (motivated by profit, that is exploitation maximizing) dominate. In short, this is a dystopia.
There is an alternative – intelligence that “recognizes the global consequences of local actions”. General guidance is given by the leadership in the form of the Golden Rule. This rule recognizes that “the collective acts as one”, that one member’s costs are another member’s revenue, and vice versa. Zero sum, win-lose, local interactions will lead to a global lose-lose. There are only win-win or lose-lose global outcomes. The problem with this conservative (viscosity-less) control is that the equilibriums of maximum activity are unstable, like an inverted pendulum, and interactions of the collective with external systems will lead to a relaxation to minimum activity. Leadership stimulates the economy of the collective to the metastable maximum, but watches helplessly as the economy crashes into economic depressions.
@@ II A collection of twinkling textures
It is instructive to look at a specific physical system that displays a rich dynamical structure. The system is an electron and an ion in a constant magnetic field. It is assumed that the magnetic field is strong enough that the electron undergoes guiding center motion, that is the electron cyclotron motion is an adiabatic invariant. It is also assumed that this magnetic field is strong enough that the motion of the electron along the magnetic field is also adiabatic. This gives two types of bound motion. The first is called Guiding Center Atoms (GCAs) (Glinsky and O’Neil, 1991), where the ion is assumed to be infinitely massive. The second is called Drifting Pairs (DPs) (Kuzmin et al., 2004), where the electron is assumed to have no mass. These two bound states are shown in Fig. 2. The motion of the GCA for two nearly identical trajectories that approach the boundary point to DP orbits is shown in Fig. 3. Note how the motion slows as the boundary point is approached, and note the difference in speed between two trajectories so that the faster trajectory laps the slower trajectory. The complete set of trajectories are shown in Fig. 4a which can be interpreted as a topographical map of the terrain shown in Fig. 4b. The motion is determined by six conserved quantities or symmetries: the energy, the action parallel to the magnetic field, the electron cyclotron action, and three total momentums.
@@
This system has two stable equilibriums, shown as the o-points, and two unstable equilibriums, shown as the x-points. One set is for the electron at 00 and $-0.3$, and another set is for the ion at $-0.9$ and $\infty$. There are basins for the GCA and the DP bound motions separated from the free motion by the black boundary called the separatrix. The o-points are stable local minimums in the energy, and the x-points are unstable saddle points that are local maximums in the energy in the vertical direction. It will take an infinite amount of time to approach the saddle points so that they will be metastable. An external thermal force coming from a heat bath will exert a force perpendicular to the motion towards lower energy. This will cause the motion to descend from the mountain top, spiralling around the mountain until the mountain pass, that is the saddle point, is reached. It will take a long time to reach the mountain pass, and an equally long time to move away from the mountain pass. The motion will then fall into one of the two basins of bound motion and spiral down to the valley center, that is the stable equilibrium. The motion will eventually wonder around the valley center in thermal equilibrium, as shown in Fig. 4c.
@@
The equilibriums shown by the o-points and the x-points are the $\beta^{*}$ singularities giving the true invariants of the motion. These equilibriums are very fundamental. They are called ground states (Weinberg, 2005), force-free states, Taylor relaxed states (Taylor, 1986), BGK modes (Bernstein et al., 1957), emergent behaviors (Gros, 2015), or most fundamentally the homology classes of the topology (Frankel, 2012).
@@ IV The underlying analytic theory
associated with the complex Lie group $\mathscr{H}=H\otimes\text{Ad}(H)$. The HJB equation is very difficult to solve analytically, but it is very easy to approximate the solution with a Multi-Layer Perceptron (MLP) (Goodfellow et al., 2016) with ReLU activation. The MLP/ReLU is a piecewise linear universal function approximator that is well suited to approximating an analytic function that is maximally flat, a solution to Laplace’s equation, except for a few isolated singularities that will have a discontinuity in the first derivative at those singularities. A block diagram of how to solve for the solution of the HJB equation using several MLPs, which is effectively a decoder into a $(P,E;Q,\tau)$ Reduced Order Model (ROM), followed by the propagator, then a matched encoder to $(p,q)$, is shown in Fig. 7.
@@ V History of social and economic collectives
As is shown in Fig. 8, US society was freed in the 1960’s. Jimmy Carter was a very well intentioned shepard when he was elected president in 1976. Unfortunately, he had no tools to control such a friction-less society and watched it wonder out of control with large amounts of inflation. The response of the political right was to reimpose large amounts of friction on US society. The implementation plan was put forward in the 1971 manifesto of Lewis Powell (Andersen, 2021), put into action by the cabal of the Koch brothers, Olin and Coors, rolled out with the election of Ronald Reagan in 1980, and is reaching its natural endpoint with the fascism of Donald Trump.
@@
Figure 8: Illustrations of different types of control. (a) Jimmy Carter tries to shepherd the new free society emerging from the 1960’s, which wonders out of control. (b) The dystopian response is Lewis Powell and the cabal locking up society. (c) Because of transactional equity and AI, we are at the down of a new era of sustainable economic prosperity, in an alpine valley shepherded by AI. Images generated by OpenAI’s DALL-E.
@@ VI Ethos of equilibriums
A being, whether that be silicon based or biological, always have reflexive circuitry, but may not have the Golden Rule (intelligent) circuitry of Fig. 1. The amount of the intelligent circuitry also varies greatly. Having the intelligent circuitry is not enough. The being must be educated. This education consists of being shown how to think critically using logic, and being exposed to many consequences of local actions, that is the arcs of history. For instance, understanding the global consequence of how Germans were treated locally after WWI compared to WWII.
@@ VII Discussion and Conclusions
The first is a deep deconvolution to expose the puppet master or master individual of the collective. This deconvolution collapses all the multiple reflections by the hall of mirrors of the puppet master back onto the primary image. This process goes by many names: renormalization, HST, Mayer Cluster Expansion, S-matrix expansion, $m$-body scattering cross sections, $m$-body Green’s functions, and the Wigner-Weyl transformation. It exposes how the collective is correlated, that is the textures of the collective. “The collective acts as one.” As was shown in Fig. 6, displaying the architecture of the Heisenberg Scattering Transformation (HST), the deep deconvolution is a logarithmic generating functional of a canonical flow of the field and the canonically conjugate field momentum $S_{p}[f(x)]$. It exposes the textures or the correlation spectrums of the motion $\left|\beta(z)\right>$. These are the functional Taylor expansion coefficients of the generating functional. Practically these can be viewed different ways. The first, for LLM, is that $S_{1}$ tells how letters are correlated, $S_{2}$ tells how letters are correlated into words, $S_{3}$ tells how words are correlated into sentences, $S_{4}$ tells how sentences are correlated into paragraphs, and $S_{m}$ how the chapters are correlated into a novel. The second, for the Mayer Cluster Expansion, is that $S_{1}$ tells how one body is correlated, $S_{2}$ tells how two bodies are correlated, $S_{3}$ tells how three bodies are correlated, $S_{4}$ tells how four bodies are correlated, and $S_{m}$ tells how $m$ bodies are correlated. The third, for Heisenberg’s S-matrix, is that $S_{1}$ is the one body scattering cross section, $S_{2}$ is the two body scattering cross section, $S_{3}$ is the three body scattering cross section, $S_{4}$ is the four body scattering cross section, and $S_{m}$ is the $m$ body scattering cross section.
@@
For the case of an economic system, the optimal unstable equilibrium is reached by stimulating the economy. This stimulation takes the form of investment. The correct level of investment is calculated by modeling, that is simulating, the economic system using the twinkling texture. An example of a twinkling texture is the price curve for a commodity, like methane, on which the economic system is based. The ponderomotive stabilization of the economy would be done by arbitrage trading methane.
@@ Acknowledgements.
First of all, we would like to thank Ted Frankel and Michael Freedman and for giving a deep appreciation for the primal importance of geometry and topology to the sciences, including economics. To the Santa Fe Institute and Murray Gell-Mann, for a stay where the richness of complex systems and how to approach them from a scientific perspective was learned. To the Institut des Hautes Etudes Scientifique and Stephane Mallat, for another stay where the methods of topological analysis were inspired. To Pastor Benjamin Larzelere for intriguing and useful discussions about religious philosophy, and to Alexandria Glinsky for insight into the theory of action painting. Last, but not least, to Robert Baird who gave a deep love and understanding of economics.
@@ Appendix A Simple Beauty and the Crazy Ones
Now we present an epilogue on the true Crazy Ones – the ones that change things. While well educated intelligence can learn the sticky textures $\beta^{*}$ from historical observation, it can not create new sticky textures. It can not create an abstract action painting if none has yet been painted. It takes theoretical play where new sticky textures are proposed, then the ramifications of those additional sticky textures explored through play. This is how the true creative Crazy Ones innovate and change things. The creative art or true genius is the choice of the new sticky textures, and the creative process is the theoretical play. So, although we are at the dawn of well educated artificial intelligence, we still must rely upon organic human creativity – the Crazy Ones!

### v6

@@ I Introduction
This method of collective control takes two forms. The first, and most common, is via embedding a viscous local force on the collective. This takes the form of rules and regulations, where penalties, both financial (fines) and physical (“lock them up”), are assessed for undesirable actions. A more subtle form of viscosity was developed during the 14th and 15th century renaissance in Northern Italy lead by the Strozzi’s and Medici’s, sanctioned by the Roman Catholic Church, and implemented by Jewish pawn brokers. This financial servitude is usury or borrow/loan debt based financing (Parks, 2005). In general, economic viscosity is debt, slavery, servitude, intolerance, and lack of freedom. In modern society, there are more nuanced implementations as interest, dividends, stock buybacks, quick Venture Capital paybacks, non-compete agreements, all-efforts clauses in employment agreements, patents, copyrights, and other labor and intellectual property restrictions. This is unintelligent or stupid control, where “the global consequences of the local actions” are not considered. The problem with this method of control has been that it has led to a degradation of the collective’s potential activity and an exploitative minimization of activity where fascism and capitalistic economic monopolies (motivated by profit, that is exploitation maximizing) dominate. In short, this is a dystopia.
There is an alternative – intelligence that “recognizes the global consequences of local actions”. It is playing the break in golf, while putting. The topography of the green is recognized, that is the curvature of the golf ball’s trajectory. This is equivalent to the underlying manifold having topology that leads to curved geodesic trajectories. General guidance is given by the leadership in the form of the Golden Rule. This rule recognizes that “the collective acts as one”, that one member’s costs are another member’s revenue, and vice versa. Zero sum, win-lose, local interactions will lead to a global lose-lose. There are only win-win or lose-lose global outcomes. The problem with this conservative (viscosity-less) control is that the equilibriums of maximum activity are unstable, like an inverted pendulum, and interactions of the collective with external systems will lead to a relaxation to minimum activity. Leadership stimulates the economy of the collective to the metastable maximum, but watches helplessly as the economy crashes into economic depressions.
@@ II A collection of twinkling textures
It is instructive to look at a specific physical system that displays a rich dynamical structure. The system is an electron and an ion in a constant magnetic field. It is assumed that the magnetic field is strong enough that the electron undergoes guiding center motion, that is the electron cyclotron motion is an adiabatic invariant. It is also assumed that this magnetic field is strong enough that the motion of the electron along the magnetic field is also adiabatic. This gives two types of bound motion. The first is called Guiding Center Atoms (GCAs) (Glinsky and O’Neil, 1991), where the ion is assumed to be infinitely massive. The second is called Drifting Pairs (DPs) (Kuzmin et al., 2004), where the electron is assumed to have no mass. These two bound states are shown in Fig. 2. The motion of the GCA for two nearly identical trajectories that approach the boundary point to DP orbits is shown in Fig. 3. Note how the motion slows as the boundary point is approached, and note the difference in speed between the two trajectories so that the faster trajectory laps the slower trajectory. The complete set of trajectories are shown in Fig. 4a which can be interpreted as a topographical map of the terrain shown in Fig. 4b. The motion is determined by six conserved quantities or symmetries: the energy, the action parallel to the magnetic field, the electron cyclotron action, and three total momentums.
@@
This system has two stable equilibriums, shown as the o-points, and two unstable equilibriums, shown as the x-points. One set is for the electron at 00 and $-0.3$, and another set is for the ion at $-0.9$ and $\infty$. There are basins for the GCA and the DP bound motions separated from the free motion by the black boundary called the separatrix. The o-points are stable local minimums in the energy, and the x-points are unstable saddle points that are local maximums in the energy in the vertical direction. It will take an infinite amount of time to approach the saddle points so that they will be metastable. An external thermal force coming from a heat bath will exert a force perpendicular to the motion towards lower energy. This will cause the motion to descend from the mountain top, spiralling around the mountain until the mountain pass, that is the saddle point, is reached. It will take a long time to reach the mountain pass, and an equally long time to move away from the mountain pass. The motion will then fall into one of the two basins of bound motion and spiral down to the valley center, that is the stable equilibrium. The motion will eventually wander around the valley center in thermal equilibrium, as shown in Fig. 4c.
@@
The equilibriums shown by the o-points and the x-points are the $\beta^{*}$ singularities giving the true invariants of the motion. These equilibriums are very fundamental. They are called ground states (Weinberg, 2005), force-free states, Taylor relaxed states (Taylor, 1986), BGK modes (Bernstein et al., 1957), emergent behaviors (Gros, 2015), or most fundamentally the homology classes of the topology (Frankel, 2012). They are force-free states because the collective system is free from the influence of external systems, that is external forces, due to the effective infinite mass of the collective system at these points since its frequency is zero there.
@@ IV The underlying analytic theory
associated with the complex Lie group $\mathscr{H}=H\otimes\text{Ad}(H)$, also known as the Weyl-Heisenberg group. The HJB equation is very difficult to solve analytically, but it is very easy to approximate the solution with a Multi-Layer Perceptron (MLP) (Goodfellow et al., 2016) with ReLU activation. The MLP/ReLU is a piecewise linear universal function approximator that is well suited to approximating an analytic function that is maximally flat, a solution to Laplace’s equation, except for a few isolated singularities that will have a discontinuity in the first derivative at those singularities. A block diagram of how to solve for the solution of the HJB equation using several MLPs, which is effectively a decoder into a $(P,E;Q,\tau)$ Reduced Order Model (ROM), followed by the propagator, then a matched encoder to $(p,q)$, is shown in Fig. 7.
@@ V History of social and economic collectives
As is shown in Fig. 8, US society was freed in the 1960’s. Jimmy Carter was a very well intentioned shepherd when he was elected president in 1976. Unfortunately, he had no tools to control such a friction-less society and watched it wander out of control with large amounts of inflation. The response of the political right was to reimpose large amounts of friction on US society. The implementation plan was put forward in the 1971 manifesto of Lewis Powell (Andersen, 2021), put into action by the cabal of the Koch brothers, Olin and Coors, rolled out with the election of Ronald Reagan in 1980, and is reaching its natural endpoint with the fascism of Donald Trump.
@@
Figure 8: Illustrations of different types of control. (a) Jimmy Carter tries to shepherd the new free society emerging from the 1960’s, which wanders out of control. (b) The dystopian response is Lewis Powell and the cabal locking up society. (c) Because of transactional equity and AI, we are at the dawn of a new era of sustainable economic prosperity, in an alpine valley shepherded by AI. Images generated by OpenAI’s DALL-E.
@@ VI Ethos of equilibriums
A being, whether that be silicon based or biological, always have reflexive circuitry, but may not have the Golden Rule (intelligent) circuitry of Fig. 1. The amount of the intelligent circuitry also varies greatly. Having the intelligent circuitry is not enough. The being must be educated. This education consists of being shown how to think critically using logic, and being exposed to many consequences of local actions, that is the arcs of history. For instance, understanding the global consequence of how Germans were treated locally after WWI compared to WWII, well educated intelligence would always enact a Marshall Plan.
@@ VII Discussion and Conclusions
The first is a deep deconvolution to expose the puppet master or master individual of the collective. This deconvolution collapses all the multiple reflections by the hall of mirrors of the puppet master back onto the primary image. This process goes by many names: renormalization, HST, Mayer Cluster Expansion, S-matrix expansion, $m$-body scattering cross sections, $m$-body Green’s functions, and the Wigner-Weyl transformation. It exposes how the collective is correlated, that is the textures of the collective. “The collective acts as one.” As was shown in Fig. 6, displaying the architecture of the Heisenberg Scattering Transformation (HST), the deep deconvolution is a logarithmic generating functional of a canonical flow of the field and the canonically conjugate field momentum $S_{p}[f(x)]$. It exposes the textures or the correlation spectrums of the motion $\left|\beta(z)\right>$. These are the functional Taylor expansion coefficients of the generating functional. Practically these can be viewed different ways. The first, for LLM, is that $S_{1}$ gives the frequency of letters, $S_{2}$ tells how letters are correlated into words, $S_{3}$ tells how words are correlated into sentences, $S_{4}$ tells how sentences are correlated into paragraphs, and $S_{m}$ how the chapters are correlated into a novel. The second, for the Mayer Cluster Expansion, is that $S_{1}$ tells how one body is correlated, $S_{2}$ tells how two bodies are correlated, $S_{3}$ tells how three bodies are correlated, $S_{4}$ tells how four bodies are correlated, and $S_{m}$ tells how $m$ bodies are correlated. The third, for Heisenberg’s S-matrix, is that $S_{1}$ is the one body scattering cross section, $S_{2}$ is the two body scattering cross section, $S_{3}$ is the three body scattering cross section, $S_{4}$ is the four body scattering cross section, and $S_{m}$ is the $m$ body scattering cross section.
@@
For the case of an economic system, the optimal unstable equilibrium is reached by stimulating the economy. This stimulation takes the form of investment. The correct level of investment is calculated by modeling, that is simulating, the economic system using the twinkling textures. An example of a twinkling texture is the price curve for a commodity, like methane, on which the economic system is based. The ponderomotive stabilization of the economy would be done by arbitrage trading methane.
@@ Acknowledgements.
First of all, we would like to thank Ted Frankel and Michael Freedman and for giving a deep appreciation for the primal importance of geometry and topology to the sciences, including economics. To the Santa Fe Institute and Murray Gell-Mann, for a stay where the richness of complex systems and how to approach them from a scientific perspective was learned. To the Institut des Hautes Etudes Scientifique and Stephane Mallat, for another stay where the methods of topological analysis were inspired. To Pastor Benjamin Larzelere for intriguing and useful discussions about religious philosophy, to Alexandria Glinsky for insight into the theory of action painting, and to Sharon Sievert for discussions on financial theory. Last, but not least, to Robert Baird who gave a deep love and understanding of economics.
@@ Appendix A Simple Beauty and the Crazy Ones
Now we present an epilogue on the true Crazy Ones – the ones that change things. While well educated intelligence can learn the sticky textures $\beta^{*}$ from historical observation, it can not create new sticky textures. It can not create an abstract action painting if none has yet been painted. It takes theoretical play where new sticky textures are proposed, then the ramifications of those additional sticky textures explored through play. This is how the true creative Crazy Ones innovate and change things. They “think different”, the title of the famous 1997 Apple ad campaign that talked about the Crazy Ones who change things and move the human race forward. The creative art or true genius is the choice of the new sticky textures, and the creative process is the theoretical play. So, although we are at the dawn of well educated artificial intelligence, we still must rely upon organic human creativity – the Crazy Ones!